# Resource-constrained Fairness

### v1

@@ 3.3.1 Performance metrics
We want to evaluate the performance of the default allocation and the fair allocation over a range of selection rates (where the number of selected instances is the same for both allocations). When the selection rate of a machine learning model is fixed, it implies that a set proportion of the instances will be chosen based on the highest scores predicted by the model, regardless of their actual scores. A useful performance metric to measure is the proportion of relevant instances among the top $R$ instances selected by the model (where $R$ denotes the resource level). This means the number of selected instances that actually have a positive target label.
@@
While precision at the top $R$ relates to the efficient use of limited resources, knowing this value also determines the values of both recall at $R$ and accuracy at $R$ for a specific dataset and a fixed value of $R$ (Rodolfa et al., 2021). We measure the cost of fairness as the loss in precision when using the fair allocation instead of the optimal allocation. For the sake of completeness, we also report the cost as loss in recall (which measures the proportion of actual positives in the top $R$ over the total number of actual positives) in Table 2 and Figure 6.
Besides the binary predictor $\hat{Y}$, each classifier will also output a prediction score S, with the interpretation that higher values of S corresponds to a greater likelihood of Y = 1 (Hardt et al., 2016). This score is transformed in a binary prediction by using a threshold: $\hat{Y}=\mathds{1}\{S>t\}$. As we use the classifiers at various thresholds, it makes more sense to evaluate the overall performance of the classifier by using the prediction scores, as this is unaffected by the choice of threshold. This is measured by the Area Under the ROC Curve (AUC). The formula for the AUC score:
$P(S(x_{i})>S(x_{z})|y_{i}=1,y_{z}=0)$
This formula measures the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance.
@@ 3.3.2 Fairness metrics
3.4 Experimental set-up
Figure 1: Experimental framework
Figure 1 illustrates our experimental framework. We use a standard train-test split, where the model is trained on the training set and the predictive performance and cost of fairness is evaluated on a separate test set.
As discussed, machine learning models do not only output prediction labels but also prediction scores. We use these scores to generate the prediction labels for a range of selection rates. For the ‘default’ labels, we select the top $R$ of instances based on the highest prediction scores, where $R$ is the available resource level, and $r$ the resulting selection rate. For the ‘fair’ labels, we compute the proportion of resources to be allocated to each group. Allocating resources for demographic parity is relatively straightforward, as we select the same percentage in each group (illustrated in Figure 2). For equal opportunity, it involves calculating the threshold needed to equalize the number of true positives. [^2] Note that the thresholds to enforce demographic parity can be calculated on the test set, but the thresholds for equal opportunity should be calculated on a separate validation set, as they require access to the target label. To ensure that the number of instances predicted as positive is equal to the available resource level, we do not calculate the thresholds but the proportions. We calculate the proportion of resources to allocate to each group on the validation set, and then enforce this on the test set for the resource level $R$. This set-up belongs to the class of post-hoc bias mitigation strategies, that attempt to make the output of machine learning model fair after the model has been trained (Mehrabi et al., 2021). We conduct this analysis across a complete range of selection rates $r$ (from 1% to 100 %) to represent different resource levels $R$.
Figure 2: Example of implementing demographic parity in a resource-constrained set-up. If we want to satisfy demographic parity for a given resource level R, we need to switch some of the selected instances from the privileged group for instances of the protected group.
The main objective of our work is to quantify the “cost of fairness", defined as the difference in performance (measured as either precision or recall) between the default and fair labeling approaches, for a fixed number of selected instances ($R$). We vary the level of available resources R, to investigate how this influences the cost of fairness.
4 Results
4.1 What influences the cost of fairness?
Table 2: Model performance (entire population, privileged group and protected group) and the average cost of fairness (measured by the loss in precision and recall) when enforcing demographic parity and equal opportunity on all datasets. The average is calculated over all the selection rates from $1\%$ to $100\%$.
@@ 4.1 What influences the cost of fairness?
In Table 2, we analyze the AUC performance of our machine learning model for different groups: the entire population, and separately for protected and privileged groups. Additionally, we report the average cost of enforcing fairness as the loss in both precision and recall between the default allocation and the fair allocation, averaged over all selection rates.
We see that the average cost of enforcing fairness is the lowest for the Fitzpatrick17K dataset, which also has a very low base rate disparity. The average cost is also very low for the Law dataset, despite its significant base rate disparities. A possible reason for this could be the relatively small size of the protected group in this dataset, which results in less reallocation of resources. The average cost of enforcing fairness across other datasets is more similar, however it is worth noting that the cost associated with enforcing both fairness metrics on CelebA is higher than for the other datasets. We see that the average loss in precision and in recall of all datasets is in line with each other, although the exact numbers differ.
It is hard to attribute the difference in costs between the datasets to a single factor, as many of the parameters will be different. This is why we perform a separate analysis on the Adult Income dataset. Here, we systematically vary key parameters—one at a time—to observe their effects on the cost of fairness. Specifically, we explore modifications to the base rate disparity, the size of the protected group, the noisiness of the whole dataset and the noisiness of the protected group. The impacts of these changes are visualized in Figure 3, with the original dataset’s results represented by a black line in each figure.
@@
Figure 4: Cost of fairness (loss in precision) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on available resource level and thus the used selection rate.
4.1.1 Base rate disparity
@@
First, we alter the base rate disparity. We see in Figures 3(a) and 3(b), that a reduction in the base rate disparity leads to a lower cost of fairness. This trend is more pronounced when enforcing DP compared to EO, but it is evident under both fairness metrics. Menon and Williamson (2018) also find that the trade-off between accuracy and fairness is determined by the strength of the correlation between the sensitive attribute and the target variable.
4.1.2 Noise
@@
We alter the performance of the model by introducing random noise to the feature values of the whole dataset. The base rates of both groups remain the same, but the model will have more trouble with correctly classifying all individuals, which results in a lower AUC score. We add noise with varying degrees from 0 to 0.5 as depicted in Figures 3(c) and 3(d). We see that when the model performance goes down, the average cost of fairness also goes down, both for DP and EO. This decrease shrinks the gap between the performances of each group, and hence reduces the overall cost of fairness. This phenomenon might also explain why the average cost of fairness is so high for the CelebA dataset in Table 2. The model is very good in distinguishing negatives and positives from each other (as measured by the AUC), so the average cost will be high. [^3] Following this reasoning, we see that if we would fit a perfect model (so the model predicts the target label perfectly for all instances) to a dataset with some level of base rate disparity, then the maximum cost would be reached at a point between the base rates and be exactly equal to the level of base rate disparity in that dataset. In the case of CelebA, the model is not perfect but the maximum cost is also relatively close to the level of the base rate disparity.
4.1.3 Subgroup noise
@@
We can also add noise only to the members of the protected group. This will lower the performance of the model for individuals of the protected group, but not for individuals of the privileged group. As shown in Figures 3(e) and 3(f), this modification increases the cost of fairness for both DP and EO, with a larger effect for EO. This effect is also larger at lower selection rates. Chen et al. (2018) also find that when there is a difference in noise level, and available covariates are not equally predictive of the outcome in both groups, fairness cannot be satisfied without sacrificing accuracy. Dutta et al. (2020) confirm that if there is not enough separability information for one group compared to the other, being fair will reduce the accuracy.
4.1.4 Size of the protected group
@@
Lastly, the size of the protected group can significantly influence the cost of fairness. When the protected group is smaller, fewer adjustments are necessary to achieve fairness, thereby reducing the cost. This relationship is clearly illustrated in Figures 3(g) and 3(h), where reducing the size of the protected group, achieved by randomly excluding part of this group, consistently lowers the cost of fairness. This is a logical finding, however, we did not see it being discussed in literature before.
4.2 What is the impact of the available level of resources (or selection rate)?
@@
Figures 4(a)-4(f) demonstrate the loss in precision for various selection rates when using a fair allocation compared to the default allocation of the machine learning model. We see that this cost varies substantially depending on the level of available resources. For the Adult (Figure 4(a)), Compas (Figure 4(b)), CelebA (Figure 4(e)) and Fitzpatrick 17K (Figure 4(f)) datasets, the cost is highest for low selection rates, while for the Dutch (Figure 4(c)) dataset, the cost is the highest for medium selection rates, and for the Law (Figure 4(d)) dataset, the cost is the highest for high selection rates. The influence of resource level on the cost of fairness has not been discussed before, yet it is a critical factor. Depending on the resource level, the cost can either be negligible or very high. The loss in recall can be seen in Figure 6. We see that the patterns are in line with the loss in precision. We can also partially agree with Rodolfa et al. (2021) who posited that for very constrained top-$k$ settings, so when the selection rate is very low, the cost of fairness can be negligible. However, we only find this to be true when the base rate is a lot higher and the performance of the model is good enough. When the base rate is low, the cost of fairness will usually be high for low-resource settings.
@@
Figure 5: Allocation results for different resource levels (Adult Income dataset). Left point on the axis represents the situation where the resources are maximally allocated to the privileged group, while the right point represents the situation where the resources are maximally allocated to the protected group.
@@
We can get more insights into how these trade-offs behave for each resource level by analyzing the different scenarios. For the Adult dataset, this is shown in Figures 5(a)-5(d). These figures demonstrate how different allocations lead to different levels of precision for one fixed resource level. The two extreme points of the curve represent all the resources being awarded to either the privileged group (red dot on the left) or the protected group (purple dot on the right). However, also for these extreme points, when the total level of resources is higher than the size of one of the groups, some of the resources will still be awarded to the other group (as can for example be seen in Figure 5(c)). We see that the optimum is not reached by awarding all the resources to one of the groups, but somewhere in the middle (darkblue dot). We note that the precision of the unconstrained (‘unfair’) model, will be very close to the optimal allocation, and that the allocations required by the fairness metrics (both DP and EO) will lead to a lower precision for every resource level (Figures 5(a)-5(d)). However, this difference is a lot larger for low resource levels (Figures 5(a)-5(b)). In Figure 5(c), the default allocation of the machine learning model is very unfair, but using the fair allocation (both DP and EO) results in a very low difference in precision. This demonstrates that although the initial model allocation is very unfair, this does not necessarily result in a high cost of fairness. In Figure 5(d), the default allocation of the machine learning model is already approximately fair, and hence the cost of fairness is also low.
4.3 What is the impact of the chosen fairness metric?
@@
We can see in Table 2, and in Figures 4 and 6 that the cost of fairness is usually lower when enforcing equal opportunity than when enforcing demographic parity. This is because the number of instances that needs to be ‘swapped’ is typically lower when enforcing equal opportunity. Liu et al. (2018) also find that the selection rate enforced by equal opportunity is likely to be much closer to the optimal than the selection rate enforced by demographic parity, while Hardt et al. (2016) show that enforcing equal opportunity leads to less loss in profits than enforcing demographic parity.
@@
The setting where EO would results in a higher number of swapped instances, is when an equal true positive rate results in stricter requirements than an equal positive rate. A case where this could happen is when one of the groups has significantly lower data quality. Demographic parity suffers less under this scenario, as here we do not care if the predictions are accurate or not. However, to enforce equal opportunity, the algorithm must do equally bad on all groups, leading to a worse overall performance. Indeed, as shown in Figures 3(e) and 3(f), the cost of enforcing Equal Opportunity (EO) increases significantly more rapidly compared to the cost of enforcing Demographic Parity (DP) when one subgroup becomes noisier.
It is important to note that in our implementation, we enforce fairness perfectly. Many papers study a relaxed formulation of fairness constraints, where some level of unfairness is allowed (Liu et al., 2018; Friedler et al., 2019; Haas, 2019).
4.4 Summarizing the costs
We can approximate the overall costs as follows: the cost of fairness depends strongly on two factors (i): the size of the disadvantaged group (if they are small, treating them differently will have little effect on global measures). (ii) The performance gap between the subsets of data that are treated differently under fairness (i.e. those individuals in the protected group that gain an advantage because of the use of fairness, and those that become disadvantaged in the privileged group due to fairness). This second factor explains much of what we see. For example, the increase of global noise decreases performance for both groups, at broadly similar proportions. This overall decrease shrinks the gap between the per-group performance, reducing the overall cost of fairness. Similarly, increasing classification noise for the worst performing group or increasing the difference in base rates increases the cost of fairness.
@@ 4.4 Summarizing the costs
Another key finding is that the cost is typically bell-shaped with respect to the global selection-rate. The highest costs occur when the selection rates are closest to the selection rates of an unconstrained classifier. This means that many of the existing works on the cost of fairness (Menon and Williamson, 2018; Friedler et al., 2019; Hort et al., 2023) will overestimate the cost of fairness for resource constrained scenarios. For scenarios where the selection rate substantially exceeds the base-rate, this is because for any informative classifier, the majority of candidates that might be positive would be selected early on, regardless of which group they belong to, and the remaining pool of candidates are likely to have negative labels regardless of which group they belong to. Similar behaviour occurs with low selection-rates. If a classifier is correctly confident that a small subset of each group do take a positive label then fairness can be enforced using these subsets for little cost. However, it must be emphasized that this behaviour is not mathematically guaranteed, and if the classifier response is genuinely uninformative within each group, then the cost of fairness will be constant for all selection rates that are below the proportion of individuals belonging to the advantaged group.
5 Discussion
So far, we measured the cost of fairness as the loss in precision by measuring the difference between the fair allocation and the default allocation of the machine learning model. However, this does not constitute the actual cost, as this is connected to the tangible impact of each correct or erroneous prediction. Consider the scenario where an algorithm can select 10,000 instances, with a loss in precision of 0.01 due to fairness constraints. This adjustments results in 100 fewer true positives and true negatives, consequently increasing the false positives and false negatives by the same number. In the context of healthcare screening, this translates to 100 missed cancer diagnoses, and 100 unnecessary follow-up tests.

### v3

@@ 3.2.1 Performance metrics
Each classifier returns a prediction score $S$, where higher $S$ values indicate a greater likelihood of $Y=1$. This score is converted to a binary prediction using a threshold: $\hat{Y}=\mathds{1}\{S>t\}$. Evaluating classifier performance with the prediction scores, unaffected by the threshold choice, is best measured by the Area Under the ROC Curve (AUC). The AUC score is given by:
$P(S(x_{i})>S(x_{z})\mid y_{i}=1,y_{z}=0)$
This formula measures the probability that the classifier ranks a randomly chosen positive instance higher than a randomly chosen negative instance.
We evaluate the performance of the default allocation and the fair allocation over a range of selection rates (where the number of selected instances is the same for both allocations). When the selection rate of a machine learning model is fixed, it implies that a set proportion of the instances will be chosen based on the highest scores predicted by the model, regardless of their actual scores. A useful performance metric to measure is the proportion of relevant instances among the top $R$ instances selected by the model (where $R$ denotes the resource level). This means the number of selected instances that actually have a positive target label.
@@
While precision at the top $R$ relates to the efficient use of limited resources, knowing this value also determines the values of both recall and accuracy at $R$ for a specific dataset and model and a fixed value of $R$ (Rodolfa et al., 2021). We measure the cost of fairness as the loss in precision when using the fair allocation instead of the optimal allocation. For the sake of completeness, we also report the cost as loss in recall [^5] The proportion of actual positives in the top $R$ over the total number of actual positives in Table 2 and Figure 5.
@@ 3.2.2 Fairness Metrics
4 What influences the Cost of Fairness?
Table 2: Model performance (entire population, privileged group and protected group) and the average cost of fairness (the loss in precision and recall) when enforcing DP and EO. The average is calculated over all the selection rates from $1\%$ to $100\%$.
@@ 4 What influences the Cost of Fairness?
In Table 2, we analyze the AUC performance of our machine learning model for different groups: the entire population, and separately for protected and privileged groups. Additionally, we report the average cost of enforcing fairness as the loss in both precision and recall between the default allocation and the fair allocation, averaged over all selection rates. We see that the average cost of enforcing fairness is the lowest for the Fitzpatrick17K dataset, which also has a very low base rate disparity. The average cost is also very low for the Law dataset, despite its significant base rate disparity. A possible reason for this could be the relatively small size of the protected group in this dataset, which results in less reallocation of resources. The average cost of enforcing fairness across other datasets is more similar, however it is worth noting that the cost associated with enforcing both fairness metrics on CelebA is higher than for the other datasets. It is hard to attribute the difference in costs between the datasets to a single factor, as many of the parameters will be different. This is why we perform a separate analysis on the Adult Income dataset.
Here, we systematically vary key parameters—one at a time—to observe their effects on the cost of fairness. Specifically, we explore modifications to the base rate disparity, the size of the protected group, the noisiness of the whole dataset and the noisiness of the protected group. The impacts of these changes are visualized in Figure 2, with the original dataset’s results represented by a black line in each figure.
@@
Figure 3: Cost of fairness (loss in precision) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). The cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.
Base rate disparity
@@
First, we alter the base rate disparity. We see in Figures 2(a) and 2(b), that a reduction in the base rate disparity leads to a lower cost of fairness. This trend is more pronounced when enforcing DP compared to EO, but it is evident under both fairness metrics. Menon and Williamson (2018) also find that the trade-off between accuracy and fairness is determined by the strength of the correlation between the sensitive attribute and the target.
Noise
@@
We alter the performance of the model by introducing random noise (with varying degrees from 0 to 0.5) to the feature values of the whole dataset. The base rates of both groups remain the same, but the model struggles more with correctly classifying individuals, resulting in a lower AUC score. We see that as model performance decreases, the average cost of fairness also goes down, both for DP and EO (Figures 2(c)-2(d)). This global decrease shrinks the gap between the performances of each group, reducing the overall cost of fairness. This phenomenon might also explain why the average cost of fairness is so high for the CelebA dataset in Table 2. The model is very good in distinguishing negatives and positives from each other (as measured by the AUC), so the average cost will be high. [^6] Following this reasoning, we see that if we would fit a perfect model (so the model predicts the target label perfectly for all instances) to a dataset with some level of base rate disparity, then the maximum cost would be reached at a point between the base rates and be exactly equal to the level of base rate disparity in that dataset. In the case of CelebA, the model is not perfect but the maximum cost is also relatively close to the level of the base rate disparity, and is reached at a point between the base rates.
Subgroup noise
@@
We can also add noise only to the members of the protected group. This will lower the performance of the model for individuals of the protected group, but not for individuals of the privileged group. As shown in Figures 2(e) and 2(f), this modification increases the cost of fairness for both DP and EO, with a larger effect for EO. Chen et al. (2018) also find that when there is a difference in noise level, and available covariates are not equally predictive of the outcome in both groups, fairness cannot be satisfied without sacrificing accuracy. Dutta et al. (2020) confirm that if there is not enough separability information for one group compared to the other, being fair will reduce the accuracy.
Size of the protected group
@@
Lastly, the size of the protected group substantially influences the cost of fairness. When the protected group is smaller, fewer adjustments are necessary to achieve fairness, thereby reducing the cost. This relationship is clearly illustrated in Figures 2(g) and 2(h), where reducing the size of the protected group, by randomly excluding part of this group, consistently lowers the cost of fairness. This is an expected finding, however, we have not seen it discussed previously.
4.1 The impact of selection rate
@@
Figures 3(a)-3(f) demonstrate the loss in precision for various selection rates when using a fair allocation compared to the default allocation of the machine learning model. We see that this cost varies substantially depending on the available resources. For the Adult (Figure 3(a)), Compas (Figure 3(b)), CelebA (Figure 3(e)) and Fitzpatrick 17K (Figure 3(f)) datasets, the cost is highest for low selection rates, while for the Dutch (Figure 3(c)) dataset, the cost is the highest for medium selection rates, and for the Law (Figure 3(d)) dataset, the cost is the highest for high selection rates. The influence of resource level on the cost of fairness has not been discussed before, yet it is a critical factor. Depending on the resource level, the cost can either be negligible or very high. The loss in recall can be seen in Figure 5. We see that the patterns are in line with the loss in precision. We can also partially agree with Rodolfa et al. (2021) who posited that for very constrained top-$k$ settings, so when the selection rate is very low, the cost of fairness can be negligible. However, we only find this to be true when the base rate is a lot higher and the performance of the model is good enough. When the base rate is low, the cost of fairness will usually be high for low-resource settings.
@@
Figure 4: Allocation results for different resource levels (Adult Income dataset). Left point on the x-axis represents the resources being maximally allocated to the privileged group, while the right point represents the resources being maximally allocated to the protected group.
@@
We can get more insights into how these trade-offs behave for each resource level by analyzing the different scenarios. For the Adult dataset, this is shown in Figures 4(a)-4(d). These figures demonstrate how different allocations lead to different levels of precision for one fixed resource level. The two extreme points of the curve represent all the resources being awarded to either the privileged group (red dot on the left) or the protected group (purple dot on the right).  [^7] However, also for these extreme points, when the total level of resources is higher than the size of one of the groups, some of the resources will still be awarded to the other group (as can for example be seen in Figure 4(c)). We see that the optimum is not reached by awarding all the resources to one of the groups, but somewhere in the middle (darkblue dot). We note that the precision of the unconstrained (‘unfair’) model, will be very close to the optimal allocation, and that the allocations required by the fairness metrics (both DP and EO) will lead to a lower precision for every resource level (Figures 4(a)-4(d)). However, this difference is a lot larger for low resource levels (Figures 4(a)-4(b)). In Figure 4(c), the default allocation of the machine learning model is very unfair, but using the fair allocation (both DP and EO) results in only a little change in precision. This demonstrates that a very unfair initial model allocation does not guarantee a high cost of fairness. In Figure 4(d), the default allocation of the machine learning model is already approximately fair. Hence, the cost of fairness is also low.
4.2 The Impact of the Fairness Metric
@@
We can see in Table 2, and in Figures 34 and 5 that the cost of fairness is usually lower when enforcing equal opportunity than when enforcing demographic parity. This is because the number of instances that needs to be ‘swapped’ is typically lower when enforcing equal opportunity. Liu et al. (2018) also find that the selection rate enforced by equal opportunity is likely to be much closer to the optimal than the selection rate enforced by demographic parity, while Hardt et al. (2016) show that enforcing equal opportunity leads to less loss in profits than enforcing demographic parity.
@@
The setting where EO would results in a higher number of swapped instances, is when an equal true positive rate results in stricter requirements than an equal positive rate. A case where this could happen is when one of the groups has significantly lower data quality. DP suffers less under this scenario, as here we do not care if the predictions are accurate or not. However, to enforce EO, the algorithm must do equally bad on all groups, leading to a worse overall performance. Indeed, as shown in Figures 2(e) and 2(f), the cost of enforcing EO increases significantly more rapidly compared to the cost of enforcing DP when one subgroup becomes noisier.
4.3 Bounding the Cost of Fairness
We now formalize these costs and consider the cost of the change induced by enforcing fairness at a particular selection rate.
As we are modeling the cost of inducing fairness under rate constraints, we consider the changes in the labeling induced by swaps, where the label assigned to one instance changes from positive to negative, and the label assigned to another instance changes from negative to positive. A sequence of swaps must always preserve the selection rate, and any pair of labellings with the same selection rate, can be transformed from one to another by a sequence of swaps.
We write $p$ for the proportion of instances in the dataset swapped from positive to negative (or equivalently the proportion swapped from negative to positive). The overall cost of inducing fairness can be written as:
$\text{cost}=p\times\text{average cost of a swap}$ (7)
Note that $p$ is somewhere between 0 and 0.5, and for standard linear measures (e.g. accuracy, recall, sensitivity, and False Positive and Negative Rates) the second term lies somewhere between $-c^{-1}$, and $c^{-1}$. Here,
$c=\frac{\text{\#instances used in the measure}}{\text{\#instances in dataset}}$ (8)
i.e. $c=1$ for accuracy, and, writing $b$ for the base rate, $c=b$ for Recall and False Positive Rate, and $c=(1-b)$ for Specificity and False Negative Rate. As such, $c^{-1}p$, and half the average cost of a swap are upper bounds on the cost of enforcing fairness.
To state this less formally, the cost of fairness, in terms of change in accuracy, is at most the proportion of swaps ($p$); while the cost in terms of change of recall is at most the proportion of swaps divided by the base rate.
We can create successive bounds based on this identity.
Given a fixed global selection rate $r$, $p$ is bounded above by $r$ (all positive instances are swapped); and $(1-r)$ (all negative instances are swapped). Therefore, the cost of fairness is bounded above by $rc^{-1}$, and $(1-r)c^{-1}$. As such, if $r$ or $(1-r)$ is much smaller than $c$, there is little cost in enforcing fairness.
@@ 4.3 Bounding the Cost of Fairness
This matches our empirical finding that the cost is bell-shaped with respect to the global selection-rate. The highest costs occur when the selection rates are closest to the selection rates of an unconstrained classifier. This means that many of the existing works on the cost of fairness (Menon and Williamson, 2018; Friedler et al., 2019; Hort et al., 2023) overestimate the cost of fairness for resource-constrained scenarios. For scenarios where the selection rate substantially exceeds the base-rate, this is because for any informative classifier, the majority of candidates that might be positive would be selected early on, regardless of which group they belong to, and the remaining pool of candidates are likely to have negative labels regardless of which group they belong to.
Similarly, given $g$, the proportion of the dataset occupied by the smallest group, we know $p<g$, and the cost of inducing fairness is bounded above by $gc^{-1}$. Again, if $g$ is much smaller than $c$, there is little cost in inducing fairness. This result is supported empirically by our experiments (see Figures 2(g) and 2(h)).
Finally, looking at the second term (the average cost of relabelling), we find this second factor explains much of what we see. For example, increasing global noise decreases performance for both groups, at broadly similar proportions. This overall decrease shrinks the gap between the per-group performance, reducing the average cost of relabelling points. Similarly, increasing classification noise for the group with greatest classification error, or increasing the difference in base rates, increases the performance gap between groups and consequentially the cost of relabelling.
Bounding loss in precision
These bounds are much weaker for non-linear measures such as precision, if they exist. For precision, assuming a fixed global selection rate $r$, we have $c=r$. Consequentially, the low selection-rate bound $rc^{-1}=1$ is particularly meaningless. However, other bounds remain informative when there is a higher selection rate – particularly, the bounds $(1-r)c^{-1}$ and $gc^{-1}$, indicate that we should see similar patterns in the costs with respect to precision for higher selection rates. Empirically, we find this — precision exhibits similar trends to the linear measures of recall and accuracy. Reflecting the much weaker bound for low selection rates, it sometimes exhibits strong changes at low resource levels, which are not present for accuracy and recall. This can be seen in Figure 3(d).
4.4 Varying Global Selection Rates